# Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech

### v1

@@ 3.3 Fine-grained Annotation
To achieve fine-grained alignment of TTS systems, we first thoroughly investigate the characteristics of segmental errors and categorize them into two types: temporal modeling errors and semantic-phonetic alignment errors, details can be found in Appendix A.1. As illustrated in Figure 1, temporal modeling errors arise from the model’s inability to accurately capture the temporal dynamics of speech, resulting in issues such as mispronunciations, abnormal silences, and unnatural prosody. For mispronunciations and abnormal silences, we utilize Whisper-timestamped [^1] https://github.com/linto-ai/whisper-timestamped and Voice Activity Detection systems to annotate the timestamps of these errors. For unnatural prosody, we incorporate the capabilities of text LLMs and professional human annotators to precisely identify affected segments (details provided in the Appendix A.2). This allows us to obtain timestamps for problematic segments and annotate fine-grained preference labels for temporal modeling errors at the token level.
@@
Semantic-phonetic alignment errors, on the other hand, arise from the mapping between phonetic input and acoustic output. These errors often manifest as repetitions or truncations, where the TTS system either over-relies on its previous outputs or fails to encode essential phonetic information. Annotating these error segments relies on automatic speech recognition systems and ground truth transcripts with forced alignment. However, unlike temporal modeling errors, which affect only specific segments, repetition and truncation errors often cause cascading issues after the initial error segment. Therefore, we apply different annotation strategies for each error type. For temporal modeling errors, only the problematic segments are annotated as error segments. In contrast, for semantic-phonetic alignment errors, all segments following the onset of the issue are annotated as error segments. The token-level preference annotation can be described as an indicator function:
$I\left(y^{i}\right)=\begin{cases}1&\text{ if }y^{i}\text{ in the error segment}\\ 0&\text{ otherwise }\end{cases},$ (4)
where $i$ represents the $i$-th token of an utterance preference data.
3.4 Preference Optimization
@@
where $\beta$ is a parameter controlling the deviation between the TTS model $\pi_{\theta}$ and the base reference model $\pi_{\text{ref}}$. However, estimating a reward model is computationally expensive and introduces additional complexity during training. To address this, we adopt the DPO (Rafailov et al., 2024) approach to align the TTS system, eliminating the need for a reward model.
Following previous studies (Korbak et al., 2022; Go et al., 2023), the optimal solution to the KL-constrained reward maximization objective in Eq.(5) can be described as follows:
@@
Since the Bradley-Terry model depends only on the difference in rewards between two completions, the partition function $Z(\cdot)$ can be canceled, allowing the human preference probability to be expressed solely in terms of the optimal TTS model and the corresponding reference model. Finally, to achieve fine-grained optimization, we introduce the token function from Eq. (4) to compute selective training loss for both preferred and dispreferred samples:
$\mathcal{L}_{\text{FPO}}=-\mathbb{E}\Bigg{[}$ $\sum_{i}I(y^{i})\cdot\log\sigma\Big{(}\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}$
$-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\Big{)}\Bigg{]}.$ (8)
@@ 4 Experimental Setup
Table 1: Objective results on WER/CER, UTMOS, SECS, and bad case ratio with different languages. For WER/CER and bad case ratio, lower values indicate better performance, while for UTMOS and SECS, higher values are better. The best results in each category are highlighted in bold.
[…]
CN EN CN EN CN EN CN EN
GT 3.06 2.17 - 3.92 4.41 -
CosyVoice 8.26+0.0% 6.97+0.0% 0.65 0.68 3.74 3.98 31% 24%
SDPO 7.76-6.1% 6.54-6.2% 0.68 0.70 3.85 4.04 17% 13%
UNO 5.36-35.1% 4.79-31.3% 0.70 0.73 3.94 4.03 13% 12%
RIO 4.81-41.8% 4.45-36.2% 0.72 0.71 4.02 4.11 11% 9%
FPO 3.92-52.5% 3.15-54.8% 0.66 0.68 3.83 4.01 8% 5%
4.1 Dataset
@@
We chose CosyVoice (Du et al., 2024) as our backbone model, which was trained on a large-scale dataset comprising approximately 170,000 hours. For RLHF optimization, we use the training set of LibriTTS (Zen et al., 2019) and AISHELL-3 (Shi et al., 2020) corpora, already included in the base model’s training. This setup eliminates the influence of introducing additional high-quality data. Following prior research (Hu et al., 2024), we sample a pool of speech prompts and a separate pool of target transcripts to generate preference samples using pre-trained CosyVoice model for RLHF optimization. Using the proposed fine-grained sampling-annotation pipeline, we select pairs of samples with significant differences as training data for RLHF, resulting in 1,000 preference training utterances for preference optimization.
4.2 Baseline Systems
@@
In addition to using CosyVoice as the base model, we reproduce the following powerful preference optimization approaches based on CosyVoice as baseline systems for a comprehensive comparison:
@@
  • Speech-DPO (Rafailov et al., 2024): We adapt the popular DPO algorithm to align the base model with human feedback, where preference data is optimized at the utterance level, similar to the method described in (Zhang et al., 2024), and denote it as SDPO.
@@
  • UNO (Chen et al., 2024a): A recent study proposes a RLHF framework tailored for TTS optimization, specifically considering the uncertainty arising from the inherent variability in subjective human speech perception and evaluation.
@@
  • RIO (Hu et al., 2024): It proposes a reverse inference optimization method based on the Bayesian principle to select exemplars from speech samples generated by the TTS system, enabling preference optimization to enhance the robustness of the TTS system.
4.3 Evaluation Metrics
@@
Subjective Metrics: We employ the naturalness mean opinion score (NMOS) to evaluate the naturalness of the generated samples. We invited 15 participants to listen to the audio samples and rate the naturalness on a 5-point scale, where 1 indicates very unnatural and 5 indicates completely natural. Furthermore, we conduct an ABX test where participants listen to two generated speech samples from different models but based on the same input and then choose the sample that sounds more natural. If the samples are too close to distinguish, they are instructed to indicate a tie.
@@
Objective Metrics: For objective evaluation, we use the speaker embedding cosine similarity (SECS), calculated using pre-trained automatic speaker verification models [^2] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification, to evaluate speaker similarity. Word error rate (WER) and character error rate (CER), transcribed by an automatic speech recognition model [^3] https://huggingface.co/openai/whisper-medium, are employed to evaluate robustness for Mandarin and English datasets, respectively. Additionally, speech quality is measured using a predicted mean opinion score (UTMOS), computed by a neural network model [^4] https://github.com/sarulab-speech/UTMOS22. Meanwhile, following previous studies (Hu et al., 2024), we introduce the bad case ratio to evaluate the robustness of different baseline systems, where samples with UTMOS below 3 or WER exceeding 20% are classified as bad cases.
5 Experimental Results
5.1 Objective Evaluation
We first compare our proposed FPO with baseline systems using objective metrics, results are shown in Table 1. For WER and CER results, FPO achieves significant improvements in speech intelligibility that reduces CER by 52.5% (from 8.26 to 3.92) and WER by 54.8% (from 6.97 to 3.15) compared to the base model CosyVoice. This outperforms all baselines, including RIO (--41.8% CER and --36.2% WER) and UNO (--35.1% CER and --31.3% WER), demonstrating FPO’s superior ability to mitigate segmental errors such as mispronunciations and repetitions.
On the other hand, FPO demonstrates remarkable robustness in terms of the bad case ratio, reducing it to 8% (CN) and 5% (EN), outperforming all baselines by a large margin. This highlights FPO’s ability to effectively minimize segmental errors by optimizing problematic segments. These results demonstrate the efficacy of fine-grained preference optimization in addressing segmental errors and enhancing the overall reliability of TTS systems.
For metrics such as SECS and UTMOS, our proposed FPO achieves only slightly higher results than the base model, while RIO and UNO achieve better performance than FPO. This is because FPO focuses on addressing localized errors rather than global attribute alignment, such as timbre or overall quality. Since the base model already demonstrates strong speaker similarity and speech quality, improvements from baseline systems are marginal, serving as a minor enhancement. We believe this trade-off is justified, as FPO prioritizes resolving segmental errors, emphasizing improvements in intelligibility and robustness over enhancements in global attributes.
@@ 5.1 Objective Evaluation
Figure 3: Violin plots of the subjective NMOS for speech generated between baseline systems and FPO.
5.2 Subjective Evaluation
To further evaluate the effectiveness of FPO, we conducted subjective evaluations on the naturalness of the generated samples. As shown in Figure 3, the violin plots illustrate the NMOS distribution, comparing the performance of FPO with baseline systems. The distribution range of CosyVoice is wider than that of other systems, with elongated density tails indicating the presence of notable outliers. This suggests instability in the naturalness performance of the generated samples. In contrast, the distribution range of FPO is more compact than CosyVoice and SDPO, demonstrating improved robustness after preference optimization. Although UNO and RIO also achieve relatively compact NMOS distributions, FPO outperforms UNO in interquartile range and shows a better overall NMOS distribution compared to RIO. A comparative analysis of the violin plots reveals significant differences between FPO and the baseline systems, particularly in median values and distribution patterns. These results highlight the superior performance of FPO in achieving more natural speech generation.
@@ 5.2 Subjective Evaluation
Figure 4: Results for ABX preference test, while "Base" denote the original CosyVoice model.
@@
Figure 5: Results of FPO and baseline systems using different sizes of training data.

### v2

@@ III-C Fine-grained Annotation
For mispronunciations and abnormal silences, we employ Whisper-timestamped [^1] https://github.com/linto-ai/whisper-timestamped and voice activity detection systems to annotate the timestamps of these errors. To address unnatural prosody, we leverage the capabilities of text LLMs and professional human annotators to precisely identify affected segments. Specifically, we focus on abnormal pauses and annotate fine-grained prosody errors through a three-step process. First, we use an automatic speech recognition system to transcribe the speech into text. Next, we employ text LLMs to determine optimal pause placements and insert break tags accordingly. Finally, professional annotators evaluate the speech samples using these break tags, identifying preferred and dispreferred prosody patterns. Fine-grained timestamps are then obtained using force-alignment tools based on the text and break tags. This allows us to obtain timestamps for problematic segments and annotate fine-grained preference labels for temporal modeling errors at the token level.
@@
Semantic-phonetic alignment errors, on the other hand, arise from the mapping between phonetic input and acoustic output. These errors often manifest as repetitions or truncations, where the TTS system either over-relies on its previous outputs or fails to encode essential phonetic information. Alternatively, truncation errors may occur when the language model fails to encode key phonetic information from the input or neglects certain critical elements, leading to synthesized speech that lacks completeness and accuracy. Annotating these error segments relies on automatic speech recognition systems and ground truth transcripts with force alignment tools. However, unlike temporal modeling errors, which affect only specific segments, repetition and truncation errors often cause cascading issues after the initial error segment. Therefore, we apply different annotation strategies for each error type. For temporal modeling errors, only the problematic segments are annotated as error segments. In contrast, for semantic-phonetic alignment errors, all segments following the onset of the issue are annotated as error segments. The token-level preference annotation can be described as an indicator function:
[…]
where $i$ represents the $i$-th token of preference data.
In the fine-grained annotation process, preference data is first identified, followed by the annotation of fine-grained preference labels by human annotators or automated agents—steps not included in conventional annotation methods. While fine-grained annotation involves additional steps, it requires less data than the traditional approach. Although utterance-level annotation is faster for human annotators, it demands more data to achieve the same level of alignment as fine-grained preference optimization. Therefore, we believe the additional cost of fine-grained annotation is both more acceptable and more efficient than conventional utterance-level annotation.
III-D Preference Optimization
@@
where $\beta$ is a parameter controlling the deviation between the TTS model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$. However, estimating a reward model is computationally expensive and introduces additional complexity during training. To address this, we follow the DPO approach [rafailov2024dpo] to directly align the TTS system using preference data, eliminating the need for a reward model.
Following previous studies [korbak2022reinforcement, go2023aligning], the optimal solution to the KL-constrained reward maximization objective in Eq.(5) can be described as follows:
@@
Since the Bradley-Terry model depends only on the difference in rewards between two completions, the partition function $Z(\cdot)$ can be canceled, allowing the human preference probability to be expressed solely in terms of the optimal TTS model and the reference model. Finally, to achieve fine-grained optimization, we introduce the indicator function from Eq. (4) to compute selective training loss for both preferred and dispreferred samples:
$\mathcal{L}_{\text{FPO}}=-\mathbb{E}\Bigg[$ $\sum_{i}I(y^{i})\cdot\log\sigma\Big(\beta\log\frac{\pi_{\theta}(y_{w}^{i}|x)}{\pi_{\text{ref}}(y_{w}^{i}|x)}$
$-\beta\log\frac{\pi_{\theta}(y_{l}^{i}|x)}{\pi_{\text{ref}}(y_{l}^{i}|x)}\Big)\Bigg].$ (8)
@@ IV Experimental Setup
TABLE I: Objective evaluation results on WER/CER, SECS, and bad case ratio with different languages. For WER/CER and bad case ratio, lower values indicate better performance, while for SECS, higher values are better. The results are averaged by five inference times and the best in each category are highlighted in bold.
test-zh test-en test-hard UTMOS $\uparrow$ Bad Case (%) $\downarrow$ F0 Corr. $\uparrow$
CER $\downarrow$ SECS $\uparrow$ WER $\downarrow$ SECS $\uparrow$ WER $\downarrow$ SECS $\uparrow$
CosyVoice [du2024cosyvoice] 3.63-0.0% 0.723 4.29-0.0% 0.609 11.75-0.0% 0.709 3.38±0.11 21% 0.71
   + SDPO [rafailov2024dpo] 2.69-25.9% 0.731 3.41-20.5% 0.645 9.84-16.3% 0.716 3.44±0.10 17% 0.70
   + UNO [chen2024uno] 2.57-29.2% 0.735 3.06-28.7% 0.667 9.06-22.9% 0.722 3.67±0.11 15% 0.75
   + RIO [hu2024rio] 2.21-39.1% 0.744 2.89-32.6% 0.681 9.12-22.4% 0.727 3.64±0.10 12% 0.73
   + Mask-DPO [gu2025maskdpo] 2.61-28.1% 0.734 3.27-23.8% 0.652 9.52-19.0% 0.715 3.48±0.10 14% 0.65
   + FPO (Ours) 1.83-49.6% 0.728 2.77-35.4% 0.614 7.98-32.1% 0.708 3.47±0.09 9% 0.78
CosyVoice2 [du2024cosyvoice2] 1.45-0.0% 0.748 2.57-0.0% 0.652 8.08-0.0% 0.730 3.49±0.11 14% 0.75
   + SDPO [rafailov2024dpo] 1.43-1.4% 0.745 2.71+5.4% 0.673 8.32+3.0% 0.728 3.59±0.08 15% 0.73
   + UNO [chen2024uno] 1.37-5.5% 0.759 2.43-5.4% 0.669 7.78-3.7% 0.736 3.61±0.11 12% 0.75
   + RIO [hu2024rio] 1.40-3.4% 0.751 2.47-3.9% 0.675 7.94-1.7% 0.737 3.73±0.12 9% 0.77
   + Mask-DPO [gu2025maskdpo] 1.42-2.1% 0.754 2.51-2.3% 0.671 7.92-2.0% 0.733 3.64±0.10 13% 0.67
   + FPO (Ours) 1.32-9.0% 0.747 2.24-12.8% 0.654 6.92-14.4% 0.731 3.56±0.10 8% 0.81
Llasa [ye2025llasa] 1.89-0.0% 0.669 3.22-0.0% 0.572 12.13-0.0% 0.638 3.54±0.12 25% 0.73
   + SDPO [rafailov2024dpo] 1.73-8.5% 0.685 2.83-12.1% 0.599 11.23-7.4% 0.633 3.67±0.11 22% 0.75
   + UNO [chen2024uno] 1.64-13.2% 0.701 2.74-14.9% 0.623 10.84-10.6% 0.663 3.65±0.13 18% 0.74
   + RIO [hu2024rio] 1.58-16.4% 0.709 2.79-13.4% 0.625 9.45-22.1% 0.671 3.75±0.11 15% 0.78
   + Mask-DPO [gu2025maskdpo] 1.71-9.5% 0.694 2.84-11.8% 0.591 9.92-18.2% 0.637 3.64±0.10 19% 0.64
   + FPO (Ours) 1.47-22.2% 0.672 2.46-23.6% 0.571 8.42-30.6% 0.640 3.58±0.12 11% 0.79
IV-A Model Configuration
We select the open-source CosyVoice [du2024cosyvoice], CosyVoice2 [du2024cosyvoice2] and Llasa [ye2025llasa] as our backbone models. These systems are trained on large-scale datasets and demonstrate impressive performance. CosyVoice series consists of a text encoder, a speech tokenizer, a speech language model and a conditional flow matching model. Specifically, the text encoder aligns the semantic spaces of text and speech tokens, while the speech tokenizer extracts semantic tokens. The speech language model then predicts speech tokens based on text inputs and an acoustic prompt. A conditional flow matching model converts the predicted speech tokens into a mel-spectrogram, which is finally reconstructed into speech signals by a pre-trained vocoder. In contrast, Llasa bypasses flow matching and the vocoder, directly reconstructing the predicted speech tokens into waveforms using a codec decoder. Our study focuses on optimizing the speech language model using fine-grained preference data. For model training, we use a learning rate of 2e-6 and a batch size of 64, optimized with the AdamW optimizer [loshchilov2017adamw]. The model is trained for only two epochs.
IV-B Dataset
@@
To construct the training data for preference optimization, we use the LibriTTS [zen2019libritts] and AISHELL-3 [shi2020aishell3] corpora, which are already included in the base models’ training. This setup eliminates the influence of introducing additional high-quality data. Following a prior study [hu2024rio], we sample a pool of speech prompts containing 100 individual speakers and a separate pool of target transcripts to generate preference samples using the pre-trained models. With the proposed fine-grained sampling–annotation pipeline, we then select sample pairs with substantial differences to build the training set, resulting in 1,000 utterances for preference optimization.
For evaluation, we use the Seed-TTS eval dataset [anastassiou2024seedtts] to evaluate model performance after preference optimization. This test set includes samples extracted from publicly available English (EN) and Mandarin (ZH) corpora and is designed to measure performance across various objective metrics. Each sample in the English and Mandarin test sets consists of one reference utterance and one target utterance spoken by the same speaker. The optimized model generates speech for the target text using the reference speech as an audio prompt.
IV-C Baseline Systems
@@
In addition to using the CosyVoice series as base models, we reproduce the following strong RL-based speech approaches based on CosyVoice to serve as powerful baseline systems for comprehensive comparison:
@@
  • Speech-DPO [rafailov2024dpo]: We adapt the popular DPO algorithm to align the base model with human feedback, optimizing preference data at the utterance level, similar to the approach described in [zhang2024speechalign]. However, instead of using ground-truth and generated samples as pairs, we use both generated samples as the preference and the other as the dispreference. We refer to this baseline as SDPO.
@@
  • UNO [chen2024uno]: A recent study proposes an RL framework tailored for zero-shot TTS models, which integrates human feedback into the TTS learning objective and specifically considers the uncertainty arising from inherent variability in subjective human perception and evaluation. The preference optimization maximizes the utility of these samples in an uncertainty-aware manner.
@@
  • RIO [hu2024rio]: An RL-based preference optimization approach tailored to the zero-shot TTS system with neural codec modelling. It introduces a reverse inference optimization method based on the Bayesian principle to select exemplars from speech samples generated by the TTS system, which is further employed to generate the original prompt samples. By using the reverse inference method, it can effectively select preference data to enhance the capacity of an advanced zero-shot TTS system.
  • Mask-DPO [gu2025maskdpo]: A fine-grained factuality alignment method based on DPO that incorporating sentence-level factuality as mask signals and only learns from factually correct sentences in the preferred samples and prevents the penalty on factual contents in the not preferred samples, which resolves the ambiguity in the preference learning.
IV-D Evaluation Metrics
@@
Subjective Metrics: We employ the naturalness mean opinion score (NMOS) to evaluate the naturalness of the generated samples. We invited 15 participants to listen to the audio samples and rate the naturalness on a 5-point scale, where 1 indicates very unnatural and 5 indicates completely natural. We apply a t-test to assess the statistical significance between systems with 95% confidence intervals for the mean opinion scores to better reflect the variability of listener ratings. Furthermore, we conduct an ABX test in which participants listen to two generated speech samples from different models based on the same input and select the more natural-sounding sample. If the samples are too similar to distinguish, participants are instructed to indicate a tie.
@@
Objective Metrics: For objective evaluation, we use the speaker embedding cosine similarity (SECS), calculated using pre-trained WavLM-large fine-tuned automatic speaker verification models [^2] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification, to evaluate speaker similarity. For intelligibility evaluation, we employ the word error rate (WER) metric for English samples and the character error rate (CER) metric for Mandarin samples. English samples are transcribed using the Whisper-Large-v3 model [radford2022whisper] [^3] https://huggingface.co/openai/whisper-large-v3, while Mandarin samples are transcribed using the Paraformer model [gao2022paraformer] [^4] https://huggingface.co/funasr/paraformer-zh. Additionally, speech quality is further measured using a predicted mean opinion score (UTMOS), computed by a neural network model [^5] https://github.com/sarulab-speech/UTMOS22. Meanwhile, following previous studies [hu2024rio], we introduce the bad case ratio to evaluate the robustness of different baseline systems. Samples are classified as bad cases if their UTMOS is below 3, or if their WER exceeds 3% on the test-zh and test-en sets, or 6% on the test-hard set. Finally, we use the Pearson correlation (F0 Corr.) between the generated and ground-truth pitch sequences as a metric to evaluate prosody performance.
V Experimental Results
V-A Objective Evaluation
We first compare our proposed FPO with baseline systems using objective metrics, results are shown in TableI. For WER and CER, FPO consistently achieves the most significant improvements in speech intelligibility across all backbones. With the CosyVoice model, FPO reduces CER by 49.6% (from 3.63 to 1.83) and WER by 35.4% (from 4.29 to 2.77), outperforming all baseline systems, including RIO (-39.1% CER, -32.6% WER) and Mask-DPO (-28.1% CER, -23.8% WER). Similarly, when built upon the stronger CosyVoice2 backbone, FPO further achieves the lowest CER (1.32) and WER (2.24 and 6.92 on the test-en and test-hard sets, respectively), surpassing other optimization methods such as UNO and SDPO. Notably, FPO also maintains desired SECS and UTMOS scores, indicating improved naturalness without compromising perceptual quality.Beyond CosyVoice-based models, we additionally evaluate FPO on Llasa, a more recent TTS backbone. FPO again demonstrates substantial intelligibility gains, reducing CER by 22.2% (from 1.89 to 1.47) and WER by 23.6% (from 3.22 to 2.46), while outperforming all other baselines, including Mask-DPO and RIO.
In terms of robustness, FPO achieves the lowest bad case ratios across all settings—reducing them from 21% to 9% on CosyVoice, from 14% to 8% on CosyVoice2, and from 25% to 11% on Llasa, indicating a strong capability to suppress segmental degradation such as mispronunciations and unnatural repetitions. These consistent improvements across diverse architectures validate the generalizability and effectiveness of our fine-grained preference optimization strategy in enhancing both the reliability and perceptual consistency of speech synthesis systems.
While FPO demonstrates only marginal improvements over the base model in metrics like SECS and UTMOS—with RIO and UNO outperforming it in these aspects—this discrepancy stems from the distinct optimization priorities of the approaches. FPO’s core design focuses on resolving localized segmental errors through token-level selective training, whereas baseline systems like RIO and UNO prioritize utterance-level alignment for global attributes such as speaker similarity and overall quality. The base model already achieves strong performance in speaker similarity and speech quality due to its large-scale pre-training, leaving limited room for improvement in these dimensions. Consequently, the gains from RIO and UNO—though statistically significant—represent incremental refinements rather than transformative advancements. In contrast, FPO’s targeted optimization reduces critical segmental errors by 40% in CER/WER and slashes the bad case ratio by a large margin, directly addressing the most frequent and perceptually salient issues in human evaluations. This trade-off aligns with practical priorities in real-world TTS deployment, where eliminating catastrophic local errors often outweighs minor improvements in global attributes.
@@ V-A Objective Evaluation
TABLE II: Subjective NMOS results for the generated speech between baseline systems and FPO with 95% confidence interval.
CosyVoice CosyVoice2 Llasa
Base 3.64±0.11 3.81±0.07 3.74±0.12
SDPO 3.68±0.09 3.83±0.11 3.80±0.09
Mask-DPO 3.60±0.12 3.79±0.09 3.71±0.11
UNO 3.69±0.11 3.72±0.10 3.75±0.08
RIO 3.76±0.08 3.87±0.08 3.82±0.09
FPO 3.83±0.09 3.91±0.08 3.93±0.10
@@
Figure 3: Results of ABX preference test, while ”C1” and ”C2” denote CosyVoice and CosyVoice2 model, respectively. MDPO denotes Mask-DPO method.
@@
Figure 4: Results of FPO and baseline systems using different scales of training data.
V-B Subjective Evaluation
To further evaluate the effectiveness of FPO, we conduct subjective evaluations on the naturalness of the generated samples. We can find that our proposed FPO consistently achieves the highest NMOS scores among all baseline RLHF alignment approaches, as shown in Table II. While several baselines (e.g., RIO and SDPO) exhibit modest increases in naturalness, FPO provides the most substantial and stable enhancement, improving NMOS by +0.19, +0.10, and +0.19 points on CosyVoice, CosyVoice2, and Llasa, respectively. To further investigate the source of these perceptual gains, we combine the objective Pearson correlation results with the subjective evaluations and find that FPO outperforms all baseline systems on both metrics. This indicates that the improved naturalness is not merely due to better intelligibility or fewer segment-level errors. Instead, FPO more effectively aligns prosodic structure, directly contributing to the enhanced perceptual quality. These findings highlight FPO’s superior performance in generating more natural and consistent speech.